### v1

# Generative AI-enabled Wireless Communications for Robust Low-Altitude Economy Networking

Abstract
Low-Altitude Economy Networks (LAENets) have emerged as significant enablers of social activities, offering low-altitude services such as the transportation of packages, groceries, and medical supplies. Unlike traditional terrestrial networks, LAENets are characterized by control mechanisms and ever-changing operational factors, which make them more complex and susceptible to vulnerabilities. As applications of LAENet continue to expand, robustness of these systems becomes crucial. In this paper, we investigate a novel application of Generative Artificial Intelligence (GenAI) to improve the robustness of LAENets. We conduct a systematic analysis of robustness requirements for LAENets, complemented by a comprehensive review of robust Quality of Service (QoS) metrics from the wireless physical layer perspective. We then investigate existing GenAI-enabled approaches for robustness enhancement. This leads to our proposal of a novel diffusion-based optimization framework with a Mixture of Expert (MoE)-transformer actor network. In the robust beamforming case study, the proposed framework demonstrates its effectiveness by optimizing beamforming under uncertainties, achieving a more than 44% increase in the worst-case achievable secrecy rate. These findings highlight the significant potential of GenAI in strengthening LAENet robustness.
Index Terms:
Generative AI, wireless physical layer, low-altitude economy networking, robustness
I Introduction
Recently, advancements in wireless communications and manufacturing technologies have propelled the development of drones and other aerial devices, greatly expanding their applications in low-altitude airspace. Enabled by these technologies, the Low-Altitude Economy (LAE) has emerged as a significant sector focusing on social opportunities within low-altitude airspace (defined as elevations up to 1,000 meters). The LAE aims to harness low-flying equipment for various social activities, including air traffic management, transportation, and surveillance, to organize the Low-Altitude Economy Networks (LAENet). These networks are essential as they provide critical services such as communication, computation, storage, and sensing [1].
However, the unique low-altitude environment and the absence of unified management regulations have introduced several uncertainties to the design of LAENet. These uncertainties, including variable weather conditions, air traffic, user demand, and emergency situations, significantly heighten the requirement for LAENet devices to complete tasks in dynamic environments [1]. Therefore, robustness has been recognized as a Key Performance Indicator (KPI) that goes beyond efficiency (speed) and effectiveness (accuracy) for LAENet. Robustness refers to the ability to maintain the qualities of service (QoS) of LAENets despite the influence of uncertainties [2]. This recognition guides the LAENet’s evolution, enabling innovation and commercial activity. For example, the tech giant Uber’s drone delivery utilizes advanced artificial intelligence algorithms, GPS systems, and sensors to avoid obstacles under different weather conditions, ensuring maximum delivery safety. However, due to the complexity of the airspace environment, the high degrees of freedom, and the strong dynamics of aerial devices, building a robust wireless network platform for LAE applications is challenging.
Enhancing the robustness of the wireless physical layer requires addressing uncertainty as a critical issue. Zero-forcing techniques are applicable for accurate signal detection [3]. Deep learning-based joint source-channel coding can provide a reliable method for data transmission [4]. However, both numerical algorithms and traditional learning methods face several challenges, including heavy dependence on training data, and susceptibility to unknown environments, which hinder their effectiveness in handling uncertainty. Generative artificial intelligence (GenAI) is an emerging technology that works to analyze data distribution. GenAI has demonstrated strong capabilities in enhancing wireless communication, including physical and network layers, through cross-dimensional data processing and generation [5]. Based on the achievements in wireless communication, there are several potential advantages of using GenAI to eliminate uncertainty and construct robust LAENets. GenAI can generate additional signal data, which is difficult to obtain in extreme scenarios, to handle diverse interference patterns. Additionally, GenAI can generate clean data, thereby removing the noise uncertainty in signals. Moreover, GenAI can learn patterns from channel transmission data to realistically simulate fading effects and provide precise uncertainty estimates.
GenAI models with transformer architectures, such as Large Language Models (LLMs) and Deepseek [6], effectively capture long-range dependencies and contextual information. They are robust against noisy inputs and distribution shifts, making them useful for semantic communication, adaptive coding, and efficient spectrum utilization in wireless networks [7]. Mixture of Experts (MoE) architectures enhance computational efficiency by activating relevant expert pathways, improving scalability and reducing errors by adapting to changing channel conditions [8].
Motivated by these analyses, this paper aims to investigate the incorporation of GenAI into the development of robust LAENets. The main contributions are summarized as follows:
  • We provide an overview of LAENets and their applications, offering a detailed analysis of the robustness within LAENet. Furthermore, we introduce various GenAI models and their applications in wireless communications toward LAENets.
  • From the perspective of the physical layer, we explore how GenAI enhances the robustness of wireless communication. Moreover, we present several robustness metrics and discuss the benefits of implementing GenAI models to support the construction of LAENets.
  • We propose an MoE-transformer actor network integrated with the diffusion-based optimization framework in LAENets. Through a case study, we illustrate how GenAI can facilitate the development of robust beamforming for LAENets.
II The overall of Robust LAENet and GenAI
In this section, we first introduce the LAENet and its related applications. Further, we discuss the robust issues and challenges in the LAENets.
Figure 1: An overall of GenAI-enabled robust LAENet. Part A illustrates the architecture of LAENet from transportation, communication, and sensing. Part B introduces three aspects of robustness in LAENets. Part C presents a comparison between GenAI and deterministic AI. Part D demonstrates important GAI-enabled applications, including channel modeling, channel estimation, and beam alignment.
II-A Conceptions and Applications of LAENet
The LAE relies on the cooperation of edge equipment and low-altitude communication networks to facilitate activities, which covers three primary aspects (shown in Fig. 1 Part A):
  • Low-Altitude Transportation: One of the most prominent applications of low-altitude transportation is drone-based delivery. Companies including Amazon, Walmart, and Wing are pioneering the use of drones to transport packages, groceries, and medical supplies.
  • Low-Altitude Communication: LAENets can address the limitations of traditional communication systems in remote and rural areas. For instance, Telia Company has deployed drones equipped with portable mobile base stations to enable remote control of forestry machinery over 5G networks.
  • Low-Altitude Sensing: Drones are invaluable tools for monitoring environmental changes. Drone enables high-resolution mapping and data collection for tracking deforestation, glacier melting, and wildlife habitats. Their ability to access hard-to-reach areas ensures comprehensive coverage, supporting effective conservation strategies.
II-B The Significance of Robust LAENet
In contrast to traditional terrestrial networks, LAENets are distinguished by automatic control mechanisms and constantly shifting operational factors, making them more complex and vulnerable. As applications such as aerial deliveries grow, maintaining robustness becomes crucial. Robustness refers to the ability to maintain the QoS of LAENet despite the influence of uncertainties [2]. Many LAENet use cases are mission-critical, where even small security compromises can significantly impact QoS. Therefore, ensuring robust design and operation in these networks is essential, focusing on three core elements as shown in Fig. 1 Part B.
  • Safety: Safety in LAENets involves ensuring the protection of equipment, operators, and service recipients during operations with low-altitude devices. Ensuring QoS in these scenarios means that safety measures must consistently perform effectively, even when faced with challenges such as adverse weather, signal interference, or unexpected obstacles.
  • Security and Confidentiality: Security and confidentiality concern LAENet’s ability to maintain secure and reliable low-altitude communications while ensuring QoS amidst various uncertainties. This involves resisting external attacks, protecting against data breaches, and preventing unauthorized access, all without compromising the performance required for different applications.
  • Accuracy and Integrity: Accuracy and integrity encompass LAENet’s ability to transmit and execute instructions precisely and provide accurate feedback on the operational status of devices. This ensures that QoS is maintained, even with unpredictable environmental factors or uncertainties in data transmission.
Therefore, designing and operating robust LAENet is imperative to ensure reliable, secure, and high quality service across diverse applications.
II-C Robustness Considerations for LAENet
Figure 2: An illustration of robustness considerations in LAENet. Part A presents several QoS metrics. Part B introduces 4 uncertainty optimizations.
To effectively evaluate and enhance the robustness of LAENet, it is necessary to introduce robust QoS metrics in various applications. As shown in Fig. 2 Part A, the key applications and corresponding indicators are as follows:
  • Sensing Accuracy Indicators: These indicators evaluate spatial, frequency, and power domain features, enabling more reliable and resilient environmental perception to enhance the robustness, including time of arrival, channel state information, and received signal strength representing the performance of LAENets.
  • Communication Rate Indicators: The robustness of the communication rate can be improved by minimizing transmission failures and errors while maximizing data throughput across varying channel conditions, which can be measured by outage probability, bit error rate (BER), spectral efficiency, etc.
  • Computation Resource Indicators: These indicators enhance data transmission efficiency, reduce signaling overhead, and accelerate processing in dynamic network conditions, as measured by computation success rate, overload tolerance, and resource availability probability.
  • Control Stability Indicators: Control stability influences the speed, trajectory, and state of the device, significantly impacting channel quality in LAENets due to distance, fading, and shadowing. Control stability can be achieved by optimizing trajectory tracking accuracy, disturbance rejection, and the stability margin index to maintain stability in dynamic environments.
Besides robust indicators, the design and deployment of robust LAENets must account for these uncertainties to meet operational requirements. We present several optimization frameworks that offer increasing levels of robustness, including deterministic, stochastic, chance-constrained, and robust optimization [9], as shown in Fig. 2 Part B. Deterministic optimization enhances performance indicators under fixed conditions, boosting efficiency within a defined environment. In contrast, stochastic optimization considers uncertainty, optimizing expected performance across various scenarios. Chance-constrained optimization ensures LAENets meet predefined requirements with a probabilistic guarantee, managing error tolerance. Robust optimization, on the other hand, focuses on worst-case performance under uncertainty, ensuring LAENets operate safely in challenging conditions [9].
To optimize these robust indicators, discriminative AI methods typically are deterministic optimizations, which focus on learning decision boundaries. However, when uncertain influencing factors increase or fluctuate significantly, their performance tends to degrade rapidly. GenAI emphasizes modeling the underlying data distribution rather than decision boundaries, making it more effective for robust optimization.
III GenAI-enabled Wireless Physical Layer for Robust LAENet
In this section, we describe the overview of GenAI. Then, we present some GenAI-enabled wireless physical layer communication methods and show how these methods improve the robust QoS metrics for LAENets.
III-A Overview of GenAI
GenAI is a type of Artificial Intelligence (AI) that leverages unsupervised learning or self-supervised learning to generate new results that resemble the learned data, as depicted in Fig. 1 Part C. Typical generative AI models include Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Diffusion Models (DMs), LLM [5]. GenAI excels in capturing and modeling specific data distributions, which allows for more accurate and detailed data analysis. For example, GenAI can detect subtle deviations that may indicate operational safety issues, enabling accurate threat detection. Additionally, it can generate synthetic data to simulate potential threats, such as spoofing attacks or signal interference in low-altitude systems, providing an adaptation mechanism [7]. These features can address robust issues encountered by LAENets, providing a solid foundation for leveraging GenAI models to enhance the robustness of LAENets in wireless physical layer communications.
III-B Channel Modeling
Millimeter-wave (mmWave) technology is the primary choice in low-altitude communications, due to its substantial bandwidth and line-of-sight (LOS) capabilities. The design and evaluation of mmWave LAENets heavily depend on the availability of robust channel models in various low-altitude scenes. However, in aerial communications, this complexity is further compounded by factors such as UAV altitude, 3D orientation, and the height of surrounding buildings. In [10], the authors introduce a two-stage GenAI model for robust and versatile air-to-ground channel modeling. In the first stage, a fully connected neural network predicts the link state, determining whether the channel is in LOS, Non-Line of Sight (NLOS), or experiencing an outage. The second stage involves a conditional VAE that generates path parameters, including path losses, delays, and angles of arrival and departure. Through VAE-generated path parameters, air-to-ground channel modeling can be adapted to various low-altitude environments, such as LOS, NLOS, and outage conditions. With a more accurate channel model, the indicators, including path loss, packet delivery ratio, and fading stability, can be significantly improved. Compared with the Third Generation Partnership Project (3GPP) channel model, the proposed approach outperforms 3GPP models and the refitted 3GPP model to effectively capture complex statistical relationships in the channel modeling data [10]. The proposed method reduces path loss in various cities by 9dB to 18dB compared to the standard model, which aligns more closely with actual values. The results demonstrate the accuracy of the GenAI method in channel modeling in various low-altitude environments.
III-C Channel Estimation
In low-altitude transportation scenarios, Orthogonal Time Frequency Space (OTFS) can select the appropriate channel estimation technique to ensure robust communications. However, traditional estimators, including least squares (LS), struggle to learn channel parameters in high-speed scenarios. Motivated by the generalization capabilities of GenAI, the authors in [11] propose a GAN-based channel estimation model, which learns from channel data rather than relying on strict mathematical models, allowing it to generalize to different channel environments. The received signal is demodulated using an OTFS demodulator and then fed to the GAN block for channel estimation, where the GAN block is trained based on 10000 high-speed data sets with the receiver velocity varying from 100 to 500 km/h. Leveraging the GAN block, the received signal becomes less susceptible to the Doppler effect and multi-path distortion during channel propagation. This enables more accurate channel estimation, which can enhance indicators, such as BER, packet error probability, and diversity gain. Compared to traditional LS estimators, machine learning-based estimators, and deep learning-based estimators, the proposed GAN enhances BER performance by 70%, 55%, and 45%, respectively. Additionally, it improves robust probability indicator outage probability by 40%, 30%, and 20%, respectively [11].
III-D Beam Alignment
In LAENets, mmWave communication faces significant beam training overhead due to the mobility of a user equipment (UE) or a base stations (BS) in LAENets. Traditional beam alignment methods, such as exhaustive search, suffer from excessive overhead, as they lack closed-form expression. The authors in [12] propose a deep recurrent (DR-)VAE, designed to learn the dynamics of the strongest beam pair. The proposed DR-VAE extends the VAE framework to predict the probability of the strongest beam pair and beam training feedback, which then minimizes overhead and maximizes frame spectral efficiency. The VAE framework handles temporally correlated observations, which facilitates a deeper analysis of the implicit relationships between action-observation sequences. These relationships can optimize resource allocation effectively. Robust QoS metrics, such as spectral efficiency and link reliability, can be enhanced in capacity and reliability. Numerical evaluations employing 3D simulated beamforming at both the BS and UE demonstrate that integrating DR-VAE with beam alignment strategies reduces the Kullback-Leibler divergence between the ground truth and the learned model by approximately 95%. The proposed method outperforms the Baum-Welch algorithm, achieving an improvement in spectral efficiency of about 10% [12].
III-E Lessons Learned
Building on prior research, GenAI has shown significant potential in physical-layer wireless communications, especially in complex aerial scenarios. Its precise monitoring abilities enable aerial equipment to perform tasks accurately at high speeds, enhancing robust QoS metrics. GenAI employs a unique learning approach that maps latent distributions to key communication model parameters including path loss, delay, and angle measurements. This distributional perspective provides insights into correlated properties, enabling effective management of multiple parameters. However, current studies often rely on deterministic or stochastic optimizations, neglecting performance analysis under uncertainty conditions.
IV MoE-transformer Actor Network for Diffusion-based Optimization
In this section, we propose an MoE-transformer actor network to enhance the uncertainty-solving ability within the diffusion-based optimization framework.
IV-A Diffusion-based Wireless Optimization
GenAI models, particularly diffusion models combined with reinforcement learning, have demonstrated significant advantages in wireless network optimization [7]. These models inherently learn distributions, adapt to the effects of multi-scale fading in wireless channels, and exhibit a level of adaptability that traditional algorithms lack. For instance, the diffusion-based optimization framework GDM [13] employs a diffusion model-based actor to generate actions, as illustrated in Fig. 3 Part C. Through its multi-step denoising process, each step incrementally removes noise, enabling the actor to explore the entire state space and identify optimal actions. However, the denoising optimization process faces several challenges, particularly in uncertainty optimization. The stochastic nature of diffusion model generation can lead to increased variability in actions under the same state, affecting network convergence.
Figure 3: An illustration of GenAI-enabled robust beamforming for secure communication. Part A presents the secure communication architecture in LAENet. Part B introduces the uncertainties in LAENets. Part C presents the diffusion-based reinforcement learning framework. Part D illustrates the network structure of the proposed MoE-transformer actor network.
IV-B MoE-transformer Actor Network
Recently, Deepseek [6], a generative LLM based on reinforcement learning, has emerged as a leading example of LLM. It enhances model learning, generation capabilities, and convergence speed by integrating the MoE approach with the attention mechanism.
  • MoE: The MoE structure comprises a gating network and multiple expert networks. The gating network assigns scores to experts based on the input and activates them accordingly, allowing each expert to process data specially, leading to more precise inferences [8]. In uncertainty optimization, the MoE can select the most suitable expert to estimate the uncertainty of the current state.
  • Attention: The attention mechanism dynamically assigns weights to different input elements, allowing the model to focus on the most relevant information. Computing attention scores selectively enhances important features, improving the efficiency of information processing. Attention enables the model to capture uncertainty dependencies and contextual relationships, leading to more accurate predictions and representations.
Inspired by this, we propose an MoE-transformer actor network structure to enhance the diffusion-based wireless optimization framework, as shown in Fig. 3 Part D. Our approach employs MoE layers to capture uncertainties across different states and denoising steps while integrating a transformer with a multi-head attention mechanism to stabilize generation. By replacing the action network in the diffusion reinforcement learning framework, the proposed actor network can utilize the advantages of the MoE structure and attention mechanism to achieve more stable and robust action generation. Therefore, the proposed method leverages the exploration and generative strengths of diffusion models to mitigate uncertainties in LAENet, enhancing overall system robustness.
V Case Study: GenAI-enabled Robust Beamforming for Secure Communication
V-A Problem Description
3D beamforming dynamically adapts beam directions and power allocations in response to environmental variations and UAV mobility, ensuring reliable and high-quality signal transmission. Robust beamforming leverages spatial diversity and directional signal control to mitigate eavesdropping risks and enhance confidentiality in LAENets [14].
V-B Secure Beamformers
We consider a secure communication scenario where a terrestrial BS transmits a confidential signal along with artificial noise to communicate with legitimate UAVs and evade eavesdropping, as depicted in Fig. 3 Part A. The terrestrial BS employs a uniform planar array (UPA) with dimensions $N_{x}\times N_{y}$, whereas both legitimate UAVs and eavesdroppers (Eves) are equipped with uniform linear arrays (ULAs) comprising $N_{b}$ and $N_{e}$ antennas, respectively. Moreover, environmental factors, such as wind speed, will introduce estimation errors in the target UAV’s location. The covert nature of the eavesdropper further exacerbates uncertainty, while errors in channel estimation and angle-of-arrival estimation also influence beamforming accuracy, as shown in Fig. 3 Part B.
V-C Optimization Design
To ensure reliable communication while mitigating eavesdropping risks, we accomplish secure beamforming by solving an optimization problem. Our objective is to maximize the achievable secrecy rate (ASR) by optimizing the beamformers for both the confidential signal and artificial noise, accounting for various sources of uncertainty. To restrict the eavesdropper’s ability, we impose a constraint that limits the capacity of the eavesdropping channel to remain below a specific threshold $C_{eve}$. We designed a reward function with a penalty item to ensure reinforcement learning algorithms can learn the optimal beamforming strategy under the constraints. Specifically, the reward function is the ASR minus the capacity of the eavesdropping channel over the threshold $C_{eve}$.
To comprehensively evaluate the proposed algorithm’s effectiveness in enhancing robustness, we solve the secure beamforming via varying robustness levels of optimization.
  • Stochastic Optimization aims to maximize the mathematical expectation ASR for a given beamforming action under uncertainty. We incorporate the eavesdropping channel capacity as a penalty term to enforce constraints.
  • Chance-Constrained Optimization maximizes ASR while ensuring constraint satisfaction with a certain probability $P_{eve}$. When the constraint satisfaction probability is greater than $P_{eve}$, we only calculate the reward brought by ASR to maximize ASR from a probability perspective.
  • Robust Optimization focuses on guaranteeing worst-case performance under uncertainty. For a given beamforming action, we use the Monte Carlo method to approximate the worst-case with the lowest achievable reward. Then, we maximize this lower bound to enhance the robustness of the generated beamforming.
(a) Robust
(b) Normalized Inference Reward
Figure 4: Numerical results of the proposed algorithm and baselines. The server is equipped with an Ubuntu 22.04 operating system and powered by an Intel(R) Xeon(R) Silver 4410Y 12-core processor and an NVIDIA RTX A6000 GPU.
V-D Numerical Results
In the experiments, we consider a terrestrial BS with a $4\times 4$ UPA, a legitimate UAV, and an eavesdropper, each equipped with $N_{b}=N_{e}=6$ antennas. The beamforming design accounts for both position and channel uncertainties while adopting the same channel power gain settings as in [14], and $C_{eve}=3b/s/Hz$. In chance-constrained optimization, we set the constraint probability $P_{eve}=70\%$. We evaluate our method benchmarking with the SAC [15] and GDM [13].
The learning curve for robust optimization is presented in Fig. 4(a), where the red, green, and blue curves represent the MoE-Transformer, GDM, and SAC algorithms, respectively. Moreover, the dashed lines are the average ASR reward under robust optimization. Under the same number of training epochs, the convergence reward of our algorithm is the highest. For stochastic and chance-constrained optimization, our proposed algorithm achieves over 200% improvement compared to GDM. Compared to the SAC algorithm, our model exhibits an improvement of 27.93% and 19.37%, respectively. In the robust optimization, the GDM algorithm outperforms the SAC algorithm, indicating that the GenAI-enabled model enhances system robustness and improves its ability to handle uncertainties. Nonetheless, our proposed model still achieves the best performance, showing improvements of 44.33% and 58.70% over the two benchmark methods. These results further highlight how the MoE-transformer architecture significantly strengthens system robustness.
Finally, we compared the normalized inference reward of our proposed method. Across the three optimization paradigms—stochastic, chance-constrained, and robust optimization—each algorithm achieves the best performance or suboptimal under its own optimization target. However, robust optimization exhibits the weakest performance for the other two criteria, while the chance-constrained approach maintains a more balanced performance. This behavior can be attributed to the inherent fading characteristics of wireless channels, such as the Rayleigh fading model. The robustness optimization strategy is not effective during an outage, where we have to allocate infinite power for channel compensation. In contrast, chance-constrained optimization effectively mitigates extreme values in a probabilistic manner, resulting in better overall performance. Therefore, in practical scenarios, absolute robustness (i.e., optimizing for the worst-case in every situation) may not be the most effective approach. Instead, incorporating probabilistic error tolerance allows for the design of robust LAENets that better align with real-world requirements.
VI Conclusion
In this paper, we have explored the application of GenAI techniques to enhance the robustness of LAENets. Initially, we conducted a systematic analysis of the robustness requirements for LAENets, followed by a comprehensive review of robust QoS metrics from the perspective of the wireless physical layer. Subsequently, we have performed an in-depth analysis of existing GenAI-enabled approaches for robustness enhancement and their potential applications. Building on this, we have proposed a diffusion-based reinforcement framework with an MoE-transformer actor network. In the robust beamforming case study, the proposed framework has effectively learned the beam pattern with various uncertainties, achieving an improvement of approximately 19% ASR in stochastic and chance-constrained optimization, and around 44% ASR in robust optimization. These results underscore the critical role of GenAI in enhancing the robustness of LAENets and highlight the need for further exploration of its applications.
Biographies
Changyuan Zhao is a PhD candidate at Nanyang Technological University, Singapore.
Jiacheng Wang is a research fellow at Nanyang Technological University, Singapore.
Ruichen Zhang is a research fellow at Nanyang Technological University, Singapore.
Dusit Niyato is a chair professor with Nanyang Technological University, Singapore.
Geng Sun is a professor with Jilin University, China.
Hongyang Du is an assistant professor with The University of Hong Kong, Hong Kong.
Dong In Kim is a professor with Sungkyunkwan University, South Korea.
Abbas Jamalipour is a professor with University of Sydney, Australia.